exponential growth and decay formula and graph


1
Expert's answer
2021-11-17T18:06:33-0500

The only solutions of the differential equation "dy\/dx=ky" are the exponential functions


"y(x)=y(0)e^{kx}"

If "k>0," we have an exponential growth.


"y(x)=y(0)e^{kx}=y(0)e^{|k|x}"




f "k<0," we have an exponential decay.


"y(x)=y(0)e^{kx}=y(0)e^{-|k|x}"